# DEEP-UNFOLDING-BASED JOINT ACTIVITY AND DATA DETECTION FOR GRANT-FREE TRANSMISSION IN CELL-FREE WIRELESS COMMUNICATION SYSTEMS

### 20 Dec 2023 v1

@@ 3.3 Training Procedure
In DU-JAD, the FBS modules and AUD module undergo distinct training processes. For the FBS iterations (equal to the number of modules), we adopt the squared Frobenius norm of the discrepancy between $\mathbf{X}_{\text{D}}$ and $\mathbf{X}_{\text{D}}^{K+1}$ as the loss function for parameter training across $K$ FBS modules, i.e., $\operatorname{Loss}_{\text{FBS}}=\|\mathbf{X}_{\text{D}}^{K+1}-\mathbf{X}_{\text{D}}\|_{F}^{2}$. For the AUD module’s parameter training, we employ the empirical binary cross-entropy (BCE) as the loss: $\operatorname{Loss}_{\text{AUD}}=-\sum_{n=1}^{N}\xi_{n}\log(L_{n})+(1-\xi_{n})\log(1-L_{n})$.
4 Simulation Results
4.1 System Setup, Performance Metrics, and Baselines
@@
As in [1], we consider a cell-free communication system with $N=400$ UEs at a height of 1.65 m height and $P=20$ to $100$ APs at a height of 15 m, each with $M=4$ antennas, all uniformly distributed in a 500 m $\times$ 500 m area. The UE activity $\{\xi_{n}\}_{\forall n}$ follows an i.i.d. Bernoulli distribution with $P_{a}=0.2$. Active UEs transmit $R_{p}=50$ pilots, based on a complex equiangular tight frame, and $R_{D}=200$ QPSK signals, i.e., $B=\sqrt{0.5}$, over a 20 MHz bandwidth channel at 1.9 GHz. The UEs have 0.1 W transmit power and adhere to a 12 dB power control range. The system works under 8 dB shadow fading variance, 9 dB noise figure, and a 290 K noise temperature. Refer to [1] for more details.
@@
4.2 Simulation Results
Fig. 2: Active user detection performance.
Fig. 3: Data detection performance.

### 28 Feb 2024 v2

@@ III-C Training Procedure
In DU-JAD, the FBS modules and AUD module undergo distinct training processes. For the FBS iterations (equal to the number of modules), we adopt the squared Frobenius norm of the discrepancy between $\mathbf{X}_{\text{D}}$ and $\mathbf{X}_{\text{D}}^{K+1}$ as the loss function for parameter training across $K$ FBS modules, i.e.,
$\operatorname{Loss}_{\text{FBS}}=\left\|\mathbf{X}_{\text{D}}^{K+1}-\mathbf{X}_{\text{D}}\right\|_{F}^{2}.$ (24)
For the AUD module’s parameter training, we employ the empirical binary cross-entropy (BCE) as the loss:
$\operatorname{Loss}_{\text{AUD}}=-\sum_{n=1}^{N}\xi_{n}\log(L_{n})+(1-\xi_{n})\log(1-L_{n}).$ (25)
IV Simulation Results
IV-A System Setup, Performance Metrics, and Baselines
@@
As in [1], we consider a cell-free communication system with $N=400$ UEs at a height of 1.65 m height and $P=20$ to $100$ APs at a height of 15 m, each with $M=4$ antennas, all uniformly distributed in a 500 m $\times$ 500 m area. The UE activity $\{\xi_{n}\}_{\forall n}$ follows an i.i.d. Bernoulli distribution with $P_{a}=0.2$. Active UEs transmit $R_{\text{P}}=50$ pilots, based on a complex equiangular tight frame, and $R_{\text{D}}=200$ QPSK signals, i.e., $B=\sqrt{0.5}$, over a 20 MHz bandwidth channel at 1.9 GHz. The UEs have 0.1 W transmit power and adhere to a 12 dB power control range. The system works under 8 dB shadow fading variance, 9 dB noise figure, and a 290 K noise temperature. Refer to [1] for more details.
To evaluate the performance of the proposed DU-JAD, we consider the user detection error rate (UDER) and the average symbol error rate (ASER):
@@
[…]
To demonstrate the efficacy of our algorithm, we compare it to the following baselines: “joint AUD-CE via [20], then DD,” “joint AUD-CE via [21], then DD,” “joint AUD-CE-DD via [7],” and “box-constrained FBS algorithm [1],” which are abbreviated as “Baseline 1” through “Baseline 4.” In Baseline 1 - 4, active UEs are identified by evaluating whether the estimated channel energy, $\|\hat{\mathbf{h}}_{n}^{K+1}\|_{F}^{2}$, exceeds the threshold $T_{\text{AUD}}$, setting $\hat{\xi}_{n}$ to $1$ if true, and 00 otherwise. Given the estimated channel matrix and active UEs via Baseline 1 and 2, data detection first performs zero-forcing equalization followed by mapping the result to the nearest QPSK symbol. To accelerate convergence, we take the result of the Baseline 1 as the starting point for Baseline 3, Baseline 4 and our DU-JAD algorithm. In DU-JAD, the number of FBS modules is fixed to $K=10$, while Baseline 1 - 4 use $200$ iterations with a stopping tolerance of $10^{-3}$ for FBS. For comparison with DU-JAD, we also provide the results of Baseline 2 to 4 with only $10$ iterations. In the results shown next, we perform $5000$ Monte–Carlo trials.
IV-B Simulation Results
Figure 2: Active user detection performance.
Figure 3: Data detection performance.